jonathan5654:

6. A candy jar contains candies of the same size in three flavors: lime, cherry, and licorice. If the probability of randomly pulling out a lime candy is 3/8, and the probability of randomly pulling out a cherry candy is 2/8, what is the probability of randomly pulling out a licorice flavored candy?

Vocaloid:

all the probabilities must sum up to 1 so you must find 1 - probability of getting lime - probability of getting cherry

jonathan5654:

how do i do that

Vocaloid:

1 - 2/8 - 3/8 = ?

jonathan5654:

1/8

Vocaloid:

you're very close hint = 1 = 8/8 so 8/8 - 2/8 - 3/8 = ? (start by subtracting the numerators)

jonathan5654:

would it be 3/8

Vocaloid:

good so 3/8 = your solution